# Golden Touchstone: A Comprehensive Bilingual Benchmark for Evaluating Financial Large Language Models

### v1

@@ Appendix B Inference Template of Large Language Models
Appendix C Typical Case Study Analysis of Typical Financial NLP Tasks
@@
Qwen-2 The sentiment of the given financial post is Negative. The user is predicting that Apple’s stock price (\$AAPL) will miss or fail to meet expectations, and they also suggest that the market might be nearing its peak when everyone is optimistic about continuous upward movement.
@@
Qwen-2 Borrower, PER; France, LOC; solar grade silicon, ORG
@@
DJIA Predict the next movement of DJIA Adj based on the day’s top 25 news items. Top1: b’Haitians are so desperate for food…The Japanese comic version of Das Kapital… 0 GPT-4o [Decrease 0]
Qwen2 [Decrease 0]
Llama-3 [Decrease 0]
@@
FinMA Rise 1
Touchstone GPT 1
Financial Sentiment Analysis As demonstrated in Table LABEL:tab:fin_sent_analysis, financial sentiment classification is one of the simpler tasks for benchmarks in financial NLP, resulting in high performance across all models tested. General-purpose models (GPT-4o, Qwen-2, Llama-3) provide not only the answer but also a detailed analysis, despite not being specifically fine-tuned on the FiQA-SA dataset. In contrast, specialized models (FinGPT, FinMA, Touchstone-GPT) that have undergone instruction tuning deliver straightforward, direct responses, illustrating their efficiency and focus in domain-specific applications.
@@
Financial NER In this information extraction task, most models demonstrated an understanding of the task intent and adhered to the instructions, signifying that even in the era of large language models, models like Qwen-2 and Llama-3 actually outperformed GPT-4o. In particular, specialized models such as FinMA and Touchstone-GPT, with more comprehensive instruction tuning, responded accurately and succinctly, highlighting their enhanced capability and focus on domain-specific tasks.
Stock Movement Prediction The Stock Movement Prediction task is one of the most challenging tasks, as it requires models to predict the daily fluctuations of the DJIA based solely on the top 25 news items from social media. From the results in Table.3, it is evident that Llama-3 performed the best, yet it still falls short of practical utility. Even Touchstone-GPT, despite specialized instruction tuning, performed poorly. Our analysis suggests that the sentiment of news items may not reliably predict stock movements and that incorporating quantitative data is essential for achieving practical model performance. Similar conclusions were drawn from experiments with traditional machine learning methods like XGBoost. Nevertheless, aside from simpler tasks like financial sentiment analysis, we also pose challenging tasks such as stock prediction, which are closer to real-world applications, leaving more room for benchmark challenges and exploration. Multimodal fusion of news and quantitative data represents a promising future direction, and we look forward to seeing models excel in these tasks.

### v2

@@ Appendix C Inference Template of Large Language Models
Appendix D Typical Case Study Analysis of Typical Financial NLP Tasks
@@
Qwen-3 The sentiment of the given financial post is Negative. The user is predicting that Apple’s stock price (\$AAPL) will miss or fail to meet expectations, and they also suggest that the market might be nearing its peak when everyone is optimistic about continuous upward movement.
@@
Qwen-3 Borrower, PER; France, LOC; solar grade silicon, ORG
@@
CMIN-US Based on the news about the stock in the past five trading days…… NEWS1: Biden Urges CEOs to Improve U.S. Cybersecurity After Attacks… 0 GPT-4o Rise
Qwen3 Decrease
Llama-3 Decrease
@@
FinMA Rise
Touchstone GPT Rise
Financial Sentiment Analysis As demonstrated in Table 8, financial sentiment classification is one of the simpler tasks for benchmarks in financial NLP, resulting in high performance across all models tested. General-purpose models (GPT-4o, Qwen-3, Llama-3) provide not only the answer but also a detailed analysis, despite not being specifically fine-tuned on the FiQA-SA dataset. In contrast, specialized models (FinGPT, FinMA, Touchstone-GPT) that have undergone instruction tuning deliver straightforward, direct responses, illustrating their efficiency and focus in domain-specific applications.
@@
Financial NER In this information extraction task, most models demonstrated an understanding of the task intent and adhered to the instructions, signifying that even in the era of large language models, models like Qwen-3 and Llama-3 actually outperformed GPT-4o. In particular, specialized models such as FinMA and Touchstone-GPT, with more comprehensive instruction tuning, responded accurately and succinctly, highlighting their enhanced capability and focus on domain-specific tasks.
Stock Movement Prediction The Stock Movement Prediction task is one of the most challenging tasks, as it requires models to predict the daily fluctuations of the CMIN-US based solely on the 5-day news items. From the results in Table.4, it is evident that GPT-4o performed the best, yet it still falls short of practical utility. Even Touchstone-GPT, despite specialized instruction tuning, performed poorly. Our analysis suggests that the sentiment of news items may not reliably predict stock movements and that incorporating quantitative data is essential for achieving practical model performance. Similar conclusions were drawn from experiments with traditional machine learning methods like XGBoost. Nevertheless, aside from simpler tasks like financial sentiment analysis, we also pose challenging tasks such as stock prediction, which are closer to real-world applications, leaving more room for benchmark challenges and exploration. Multimodal fusion of news and quantitative data represents a promising future direction, and we look forward to seeing models excel in these tasks.
@@
Appendix E Information About Use Of Ai Assistants
In the preparation of this work, the author used AI-assisted technology (specifically, large language models such as GPT-4 and Deepseek-V3) exclusively for text refinement purposes. The AI was employed to assist in proofreading, correcting grammatical errors, and polishing linguistic expressions to improve the clarity and readability of the manuscript.